ISRAEL JOURNAL OF MATHEMATICS

TABLE OF CONTENTS, VOLUME 97, 1997

Bourgain, J. Estimates related to sumfree subsets of sets of integers  71

Chilin, V. I., P. G. Dodds and F. A. Sukochev The Kadec–Klee property in symmetric spaces of measurable operators

203

Chinak, M. The fixed point statements on Bergman type complex manifolds

293

Collet, P., S. Martínez and B. Schmitt The Pianigiani–Yorke measure for topological Markov chains

61

Dar, S. On the isotropic constant of non-symmetric convex bodies  151

Dodds, P. G. See V. I. Chilin

Donin, J. and S. Shnider Quasi-associativity and flatness criteria for quadratic algebra deformations

125

Downarowicz, T. The royal couple conceals their mutual relationship: A noncoalescent Toeplitz flow

239

Jakóbczak, P. Highly nonintegrable functions in the unit ball  175
Kenyon, R. Projecting the one-dimensional Sierpinski gasket  221

Kononenko, A. and C. B. Yue Cohomology and rigidity of Fuchsian groups

51

Kornfeld, I. and M. Lin Coboundaries of irreducible Markov operators on C(K)

189

Larsen, M. and R. Pink A connectedness criterion for ℓ-adic Galois representations

1

Lin, M. See I. Kornfeld

Martínez, S. See P. Collet

Miller, C. An extension of Hölder’s theorem on the Gamma function 183

Pink, R. See M. Larsen

Ransford, T. J. Norm inequalities for holomorphic semigroups  157

Reznikov, A. G. Fundamental group of self-dual four-manifolds with positive scalar curvature

93

Riley, D. M. and J. F. Semple Power closure and the Engel condition

281

Roichman, Y. Decomposition of the conjugacy representation of the symmetric groups 305 Rubshtein, B.-Z. Lacunary isomorphism of decreasing sequences of

measurable partitions

317

Sankey, A. D. Regular weights and the Johnson scheme  11

Schmitt, B. See P. Collet

Semple, J. F. See D. M. Riley

Shnider, S. See J. Donin

Sukochev, F. A. See V. I. Chilin

Wenzel, J. Vector-valued Walsh–Paley martingales and geometry of Banach spaces

29

Wojtaszczyk, P. Uniqueness of unconditional bases in quasi-Banach spaces with applications to Hardy spaces, II

253

Yue, C. B. Foliations with leaves of nonpositive curvature  113

Yue, C. B. See A. Kononenko

Zapletal, J. Strongly almost disjoint functions  101